### v1

# CMOB: Large-Scale Cancer Multi-Omics Benchmark with Open Datasets, Tasks, and Baselines

Abstract
Machine learning has shown great potential in the field of cancer multi-omics studies, offering incredible opportunities for advancing precision medicine. However, the challenges associated with dataset curation and task formulation pose significant hurdles, especially for researchers lacking a biomedical background. Here, we introduce the CMOB, the first large-scale cancer multi-omics benchmark integrates the TCGA platform, making data resources accessible and usable for machine learning researchers without significant preparation and expertise.To date, CMOB includes a collection of 20 cancer multi-omics datasets covering 32 cancers, accompanied by a systematic data processing pipeline. CMOB provides well-processed dataset versions to support 20 meaningful tasks in four studies, with a collection of benchmarks. We also integrate CMOB with two complementary resources and various biological tools to explore broader research avenues.All resources are open-accessible with user-friendly and compatible integration scripts that enable non-experts to easily incorporate this complementary information for various tasks. We conduct extensive experiments on selected datasets to offer recommendations on suitable machine learning baselines for specific applications. Through CMOB, we aim to facilitate algorithmic advances and hasten the development, validation, and clinical translation of machine-learning models for personalized cancer treatments. CMOB is available on GitHub (https://github.com/chenzRG/Cancer-Multi-Omics-Benchmark).
1 Introduction
Cancer is a major public health concern with increasing incidence and leading to mortality. The National Cancer Institute (NCI) reports that the high costs of cancer care have been projected to grow to \$246.6 billion by 2030 (COS, 2023). Without further investment in research and prevention, cancer will cost the global economy an accumulated Int\$25.2 trillion in healthcare costs over the next 30 years (Kreier, 2023). Formerly, cancer research remains at the histomorphologic level. Cancer was classified solely according to organs of the origin or simplistic appearance features (Schneider et al., 2005). It is impossible to precisely distinguish and treat many morphologically similar and organ-homologous cancers. Over the past 15 years, increasing molecular and tumor biology knowledge has proven that cancer is a multi-omics disease (Hasin et al., 2017; Olivier et al., 2019; Vasaikar et al., 2018). Cancer is influenced by events occurring at various omics levels, which are interconnected (Balmain et al., 2003; Sandoval and Esteller, 2012; Jones and Baylin, 2007). For example, changes in genome sequences (genomics) regulate gene expressions (transcriptomics), finally reflecting on the phenotype, such as cancer subtypes, responses to treatment, and prognosis (Schneider and Orchard, 2011; Mani et al., 2022). This shift toward a multi-omics perspective led to a new era of cancer research.
Machine learning has proven helpful for mining multi-omics data (Reel et al., 2021a; Picard et al., 2021). However, the development of effective approaches for studying cancer multi-omics (or omics) remains relatively limited. This issue is quite apparent compared with the more common single-cell omics studies, which are currently the main focus of machine learning scientists (Wang and Bodovitz, 2010). The main reason is the lack of available data sources. Single-cell omics data can be easily generated in controlled in-lab experiments, and tens of thousands of cell-level cultured samples are available for training machine learning models (Lee et al., 2020). Cancer omics data come from clinical environments with data collection restrictions such as patient privacy. Each research project typically involves only a limited number of patient samples (Menyhárt and Győrffy, 2021b). Actually, a large amount of cancer omics data still belongs to individual research groups, hindering scientists from formulating cancer multi-omics studies as solvable machine-learning tasks.
There is an increasing effort to make cancer omics data publicly available. The Cancer Genome Atlas (TCGA) (The Cancer Genome Atlas Research Network, 2013) is the most established and well-used public multi-omics data resource. TCGA has molecularly characterized patient samples from various cancer types, providing the research community over 2.5 petabytes of multi-omics data. Despite its potential, the usability of TCGA presents challenges, particularly for researchers without a biomedical background (Tomczak et al., 2015).
(1) Lacking omics-ready datasets. While TCGA offers many cancer samples, they are not specifically tailored for omics studies and lack cross-omics alignment. Omics data processing involves several complex steps, such as categorizing omics, sample transformation, uniforming, or quality control, all of which require domain-specific expertise and bioinformatics tools. Therefore, it is necessary to establish a benchmark of cancer multi-omics datasets that can be used directly, even by non-experts.
(2) Lacking task-ready datasets. The fundamental unit in TCGA is the patient, which lacks precise categorization for specific cancer multi-omics tasks. Satisfied samples are often scattered across repositories and presented in complex, machine-unreadable file formats (Kircher and Kelso, 2010). Extensive curation is required to link samples with metadata and convert them into formats suitable for model analysis. Making data resources ready to learn computationally on cancer multi-omics tasks is very important.
(3) Non-compatibility with cross-platform databases. Studying cancer multi-omics often requires integration with biological corpora to maintain biological accuracy. For example, transcriptomics data can be mapped to graph-structured gene corpora (e.g., STRING (Szklarczyk et al., 2023) or GEO (Barrett et al., 2012)) to study gene regulation networks or Protein-Protein Interactions (PPI). However, combining TCGA with interaction-focused databases requires bioinformatics expertise, manual retrieval, and scripting skills to ensure accurate data integration. Standardizing data formats and providing seamless data integration could enhance the development of interdisciplinary machine-learning approaches.
To address the challenges outlined above, we introduce CMOB, the first Cancer Multi-Omics Benchmark that integrate TCGA platform and make data resource more accessible and usable for machine learning researchers without significant preparation and expertise. To summarize, our contributions are as follows:
  1. 1.
    Omics-ready datasets. CMOB includes carefully processed 20 datasets covering 32 cancer types with aligned patient samples across four major omics types, i.e., mRNA expression, microRNA expression, DNA methylation, and copy number variations. We assemble multiple feature scale versions for each dataset to adapt to downstream tasks. We also provide complete data processing pipelines to facilitate replication and user customization.
  2. 2.
    Tasks-ready datasets and baselines. CMOB offers 20 learning tasks in four studies, including cancer subtype identification, pan-cancer classification, and TCGA omics data imputation, each with a corresponding dataset version. For all benchmark tasks, we collect various baseline models from both the biomedical community and recent machine learning advancements. Moreover, all models are reproduced, evaluated, and extended under our benchmark datasets.
  3. 3.
    Complementary resources. To date, CMOB includes three complementary resources: the cross-platform STRING corpus, clinical health records, and downstream analysis tools. We provide user-friendly and compatible integration scripts that enable non-experts to easily incorporate this complementary information for broader biological research and clinical support and help researchers conduct classic and reliable biological validations of their data analysis frameworks.
Remark: There is a growing body of literature studying cancer omics in machine learning (Withnell et al., 2021; Chen et al., 2022; Yang et al., 2023). The results typically involve self-collected datasets from TCGA Silva et al. (2016). Some researchers make their data collection and processing source code publicly available. However, using these resources can be cumbersome and requires substantial biological knowledge and computational tools like R packages Zhu et al. (2014). Moreover, much of this code is scattered across various cancer research articles in the biomedical community, complicating accessibility for machine learning researchers Cline et al. (2013); Liu et al. (2018). The lack of comprehensive, publicly available databases hinders the progress in cancer omics studies.
The remainder of the paper proceeds as follows. After reviewing related work in Section 2, Section 3 introduces the construction of CMOB. Before concluding, in Section 4, we conduct the baseline experiments on selected datasets and report the results.
Figure 1: CMOB provides (1) 20 cancer multi-omics datasets covering 32 cancers and a systematic data processing pipeline; (2) well-processed dataset versions to support 20 meaningful tasks with a collection of benchmarks; and (3) complementary resources to connect cross-platform databases and support fundamental cancer biology or clinical studies.
2 Related Work
CMOB is the first benchmark that includes a variety of available datasets and learning tasks for cancer multi-omics data analysis. In this section, we discuss how CMOB integrates with machine learning methods for cancer omics studies and compare it with benchmarks in related areas.
2.1 Relation to Machine-Learning Methods for Cancer Omics Study
Over the past ten years, studying cancer omics has become a cornerstone of modern cancer research Chakraborty et al. (2018); Berg et al. (2017). Key tasks in omics data analysis include identifying clinical phenotypes such as tumor types and subtypes, drug responses, and survival outcomes Yang et al. (2021a); Liu et al. (2010); Menyhárt and Győrffy (2021a); Li et al. (2021a); Wang et al. (2020); Chen et al. (2023). Researchers are dedicated to developing computational methods to accelerate all aspects of cancer omics development Kaur et al. (2021); Chen et al. (2022); Arjmand et al. (2022). Earlier studies use some statistical-based methods from the biomedical community Reel et al. (2021b). For example, iClusterBayes Mo et al. (2018) integrates multi-omics data using a Bayesian latent variable regression model for downstream prediction tasks; SNF Wang et al. (2014) constructs sample-sample similarity networks at omics levels for cancer clustering. Additionally, there are notable deep learning-based approaches from the machine learning community, such as Subtype-GAN Yang et al. (2021b), which employs a generative adversarial network (GAN) structure to extract features from each omics source for cancer subtype identification. Other works propose using variational autoencoders (VAE) (Withnell et al., 2021) and Vector Quantized-VAE (Chen et al., 2022, 2023) to capture a meaningful latent space for distinguishing cancer samples. Recent attempts in contrastive learning have shown the effectiveness of integrating multi-omics data into a unified feature space (Yang et al., 2023). Cancer omics data are also pivotal in related areas like gene network analysis Chou et al. (2014); Heo et al. (2021). For example, Saha et al. Saha et al. (2017) utilized gene expression data to infer transcriptome-wide gene networks, while Aibar et al. Aibar et al. (2017) developed a workflow to map gene regulatory networks in omics data and identify stable cell states.
2.2 Relation to Other Benchmarks on Biomedical and Healthcare Data
Compound chemical structure data is one of the most discussed data forms in the biochemical community. Several benchmarks have been proposed to advance representation learning for compounds and proteins. For example, MoleculeNet offers datasets for molecular modeling, while TAPE Rao et al. (2019) provides tasks aimed at protein transfer learning. More relevant to our focus, although primarily concentrated on drug discovery and development, TDC Huang et al. (2021) encompasses a broader range of data modalities relevant to therapeutics, including compounds, biomolecular interactions, and genomic sequences. Moreover, there are several review studies Watson et al. (2022); Colomé-Tatché and Theis (2018); Miao et al. (2021); Ma et al. (2020) on single-cell omics data integration methods within bioinformatics and omics research. However, cancer multi-omics data is relatively absent in the data science and machine learning communities as an important data form in biomedical and healthcare. We address this gap by providing cancer multi-omics datasets benchmark, meaningful learning tasks, and extensive evaluations, positioning CMOB as a critical component in biomedical and healthcare data benchmarks.
3 Cancer Multi-Omics Benchmark (CMOB)
CMOB has three major components: $\rm(\hskip 1.79993pti\hskip 1.79993pt)$ a collection of cancer multi-omics datasets and a systematic data processing pipeline; $\rm(\hskip 1.79993ptii\hskip 1.79993pt)$ well-processed dataset versions to support meaningful tasks with a collection of benchmarks; and $\rm(\hskip 1.79993ptiii\hskip 1.79993pt)$ complementary resources to connect cross-platform databases and support fundamental cancer biology or clinical studies. An overview can be found in Figure 1.
3.1 Overall Datasets Composition
Table 1: CMOB provides a collection of 20 multi-omics datasets covering 32 cancer types, that is, one Pan-cancer dataset, nine unlabeled cancer subtype datasets, five labeled, golden-standard subtype datasets, and five TCGA data imputation datasets.
Dataset Feature Scale Omics Feature Size Sample Size #Baselines Learning Task
mRNA miRNA Methy CNV
ACC Orignal 18034 368 19045 19525 177 10 Clustering
KIRP Orignal 18465 769 18715 19551 273 10 Clustering
KIRC Orignal 18464 352 19045 19523 314 10 Clustering
LIHC Orignal 17946 846 18714 19551 364 10 Clustering
LUAD Orignal 18310 427 19052 19551 450 10 Clustering
LUSC Orignal 18206 423 19060 19551 363 10 Clustering
PRAD Orignal 17954 759 19049 19568 450 10 Clustering
THCA Orignal 17261 375 19052 19551 291 10 Clustering
THYM Orignal 18354 1018 18716 19551 119 10 Clustering
Pan-cancer Aligned 3217 383 3139 3105 8314 10 Classification
GS-COAD Orignal 17261 375 19052 19551 260 10 Classification
GS-BRCA Orignal 18206 368 19049 19568 671 10 Classification
GS-GBM Orignal 20684 335 19034 19545 243 10 Classification
GS-LGG Orignal 18345 345 19023 19534 246 10 Classification
GS-OV Orignal 17354 244 19034 19534 284 10 Classification
Imp-COAD Top 2000 200 2000 2000 260 7 Imputation
Imp-BRCA Top 2000 200 2000 2000 671 7 Imputation
Imp-GBM Top 2000 200 2000 2000 243 7 Imputation
Imp-LGG Top 2000 200 2000 2000 246 7 Imputation
Imp-OV Top 2000 200 2000 2000 284 7 Imputation
As shown in Table 1, CMOB provides a collection of 20 multi-omics datasets including:
-) One pan-cancer dataset involves patients with 32 cancer types.
-) Nine unlabeled cancer subtype datasets include Adrenocortical Carcinoma (ACC), Kidney Renal Papillary Cell Carcinoma (KIRP), Kidney Renal Clear Cell Carcinoma (KIRC), Liver Hepatocellular Carcinoma (LIHC), Lung Adenocarcinoma (LUAD), Lung Squamous Cell Carcinoma (LUSC), Prostate Adenocarcinoma (PRAD), Thyroid Carcinoma (THCA), and Thymoma (THY).
-) Five labeled, golden-standard subtype datasets corresponding to five cancers include Colon Adenocarcinoma (GS-COAD), Breast invasive carcinoma (GS-BRCA), Glioblastoma Multiforme (GS-GBM), Brain Lower Grade Glioma (GS-LGG), and Ovarian Serous Cystadenocarcinoma (GS-OV).
-) Five TCGA data imputation datasets include corrupted omics profiles from the above well-studied cancer types involving Imp-COAD, Imp-BRCA, Imp-GBM, Imp-LGG, and Imp-OV.
-) Two complementary data resources include a collected corpus from STRING and a collection of Electronic Health Records (EHR) data for cancer samples, accompanied by interactive scripts for integration.
Multiple-Scaled Feature. Cancer multi-omics analysis always suffers from an unbalanced sample and feature size. CMOB hence provides three versions of feature scales, i.e., Original, Top, and Aligned, to support feasible analysis. Original features are extracted directly from each dataset and correspond to the complete set of features without filtering. Users can customize their datasets. Top features are identified through ANOVA St et al. (1989) statistical testing according to p-values, selecting the most significant features among samples. This approach unifies the feature size and potentially reduces the noise features. Aligned features are determined by the intersection of features present across all sub-datasets, corresponding to the shared features among different sub-datasets.
3.2 Omics-Ready Datasets and Assembly Pipeline
Motivation. TCGA was initially established to support broad cancer research. It is not specifically tailored for multi-omics studies but collects data at each cancer genome characterization workflow step. Collecting multi-omics data from TCGA presents several challenges:
  • No omics-specific data were given. Raw data are in many different formats, and other resource enrollments, e.g., diagnostic imaging and clinical records, are also included.
  • No unified omics profiles (feature resources) were given. Omics profiles collected for one sample are scattered across repositories.
  • Sample retrieval, profile alignment, and data cleaning by bioinformatics tools are necessary to convert raw data into omics profiles suitable for analysis.
CMOB has removed these barriers with the assembled pipeline below:
(1) Omics Recruitment and Resource Collection: CMOB recruits four representative omics types that are most discussed in current cancer studies, that is, mRNA expression (mRNA), miRNA expression (miRNA), DNA methylation (Methy), and copy number variations (CNV). More detailed information on these omics types can be found in the Supplementary Material Section B. The raw omics data are obtained from TCGA through the Genomic Data Commons (GDC) Data Portal Grossman et al. (2016). While samples are provided by different biologically experimental platforms, we incorporate data from as many platforms as possible to include a broader range of patient samples. For instance, for Methy omics data, we chose platforms like "Illumina Human Methylation 27", "Illumina Human Methylation 450", and "Illumina Human Methylation EPIC". We verify the integrity of the files using checksums (as specified in the MANIFEST file) to ensure no files are corrupted or missing. The collected data files are then organized by cancer type for further processing.
(2) Omics-specific Data Formatting and Profiling: Each type of omics data requires a specific series of processing steps tailored to its unique biological characteristics. Below, we outline a general pipeline that summarizes these processes:
  • Omics Identification and Categorization. Identify the type of omics data from the metadata and ensure proper categorization (e.g., "Transcriptome Profiling" for transcriptomics).
  • Platform and Experimental Protocol Verification. Identify and confirm the experimental platform and workflow type from metadata (e.g., Illumina Hi-Seq for sequencing platforms).
  • Feature Alignment and Profiling. Convert raw data into omics features based on technical details in metadata. For example, aligning scaled RSEM estimates to FPKM for platforms like Illumina Hi-Seq. Key notions can be found in guidelines like "Average methylation (beta-values) of promoters defined as 500bp upstream & 50 downstream of Transcription Start Site (TSS)," which shows the gene methylation definitions.
  • Mensurment Transformation and Adjustments. Apply necessary transformations (e.g., logarithmic for transcriptomics or median-centering for methylation) to ensure data consistency.
  • Annotation and Saving. Annotate processed data to ensure that omics data is labeled with unified gene IDs. Store annotated data in appropriate categories based on the specific tasks.
More details on the data processing pipeline can be found in the Appendix Section B.
3.3 Task-Ready Datasets and Collected Baselines
As introduced in Section 3.1, CMOB currently provides 20 learning tasks in three studies, including pan-cancer classification, cancer subtype identification, and omics data imputation, each with a corresponding dataset version, baseline methods, and evaluation metrics.
(1) Pan-cancer Classification:
Motivation.This task aims to identify the specific cancer type for each patient, enhancing early diagnostic accuracy and potentially improving treatment outcomes.
Baseline Methods. Several computational multi-omics data integration methods have been proposed for cancer identification using classical statistical machine learning and deep-based methods. Currently, we have enrolled well-used, open-sourced statistical methods, including: Similarity Network Fusion (SNF) (Wang et al., 2014) integrates omics data by iteratively refining sample similarity networks and applying spectral clustering; Neighborhood-based Multi-Omics clustering (NEMO) (Rappoport and Shamir, 2019) converts sample similarity networks to relative similarity for group comparability; Cancer Integration via Multi-kernel Learning (CIMLR) (Wilson et al., 2019) combines various Gaussian kernels into a similarity matrix for clustering; iClusterBayes (Mo et al., 2018) projects input into a low-dimensional space using the Bayesian latent variable regression model for clustering; moCluster (Meng et al., 2016) uses multiple multivariate analyses to calculate latent variables for classification; We also enrolled more recent deep-dased methods, including: Subtype-GAN (Yang et al., 2021b) extracts features from each omics data by relatively independent GAN layers and integrates them; DCAP (Chai et al., 2021) integrates multi-omics data by the denoising autoencoder to obtain the representative features; MAUI (Ronen et al., 2019) uses stacked VAE to extract many latent factors to identify patient groups; XOmiVAE (Withnell et al., 2021) uses VAE for low-dimensional latent space extraction and classification; MCluster-VAEs (Rong et al., 2022) uses VAE with an attention mechanism to model multi-omics data.
Evaluation Metrics. Referring to Golub et al. (1999), we propose precision (PREC) Hossin and Sulaiman (2015), normalized mutual information (NMI) Ferri et al. (2009), and adjusted rand index (ARI) Santos and Embrechts (2009) to evaluate the degree of agreement between the subtyping results obtained by different methods and the true labels.
(2) Cancer Subtype Clustering and Golden-Standard Subtype Classification:
Motivation.
Each specific cancer comprises multiple subtypes. Cancer clustering or classification aims to categorize patients into subgroups based on their multi-omics data. The reason is that while the subtypes may differ in their biochemical levels, they often share the same morphological traits, such as physical structure and form in an organism. However, for most cancer types, subtyping a cancer is still an open question under discussion. Thus, cancer subtyping tasks are typically clustering tasks without ground true labels. Here, the cancer research community has thoroughly analyzed the subtypes of some of the most common cancer types in a previous study. Therefore, we consider these subtypes to contain the true labels and set up a classification task for these subtypes.
Baseline Methods. Since most methods do not have a specific to apply for labeled or unlabeled datasets, they can serve as baselines across both types of tasks. We use the same baselines (i.e., SNF, NEMO, CIMLR, iClusterBayes, moCluster, Subtype-GAN, DCAP, MAUI, XOmiVAE, and MCluster-VAEs) as in pan-cancer classification tasks.
Evaluation Metrics. For subtype clustering, we evaluate the baseline results using the silhouette coefficient (SIL) Shahapure and Nicholas (2020) and log-rank test p-value on survival time (LPS) Xie and Liu (2005). For the golden-standard subtype classification, we also use the metrics of PREC, NMI, and ARI.
(3) Omics Data Imputation:
Motivation. We also set up an essential learning task focused on omics data. The collected Omics data are typically unified with several missing values due to experimental limitations, technical errors, or inherent variability. The imputation process is crucial for ensuring the integrity and usability of TCGA omics data (You et al., 2020).
Baseline Methods. There are several well-used methods for imputing missing values in datasets. Currently, we enrolled six of them, including: Mean imputation (Mean) imputes missing values using the mean of all observed values for the same feature; K-Nearest Neighbors (KNN) imputes missing values using the K-nearest neighbors with observed values in the same feature. The weights are based on the Euclidean distance to the sample; Multivariate imputation by chained equations (MICE) (van Buuren and Groothuis-Oudshoorn, 2011) runs multiple regressions where each missing value is modeled based on the observed non-missing values; Iterative SVD (SVD) (Troyanskaya et al., 2001) uses matrix completion with iterative low-rank SVD decomposition to impute missing values; Spectral regularization algorithm (Spectral) (Mazumder et al., 2010) is a matrix completion model that uses the nuclear norm as a regularizer and imputes missing values with iterative soft-thresholded SVD; Graph neural network for tabular data (GRAPE) (You et al., 2020) transforms rows and columns of tabular data into two types of nodes in the graph structure. Then, it uses a graph neural network to learn node representations and turns the imputation task into a missing edge prediction task on the graph; Generative Adversarial Imputation Nets (GAIN) (Yoon et al., 2018) imputes missing data by leveraging the adversarial process to learn the underlying distribution.
Evaluation Metrics. We use metrics including mean absolute error (MAE) and root mean squared error (RMSE), which are commonly used to assess imputation quality.
3.4 Complementary Resources
Researchers often integrate multi-omics data with complementary datasets in omics studies to explore broader research avenues. The main efforts often incorporate biological knowledge corpus to develop robust machine learning models, drawing on resources like the STRING database (Szklarczyk et al., 2023). Hence, CMOB not only curates cancer multi-omics data but also seamlessly bridges it with auxiliary datasets.
(1) The first inclusion is the STRING API, a renowned web-based tool for bio-network analysis of graph-structured data. CMOB provides a comprehensive Mapping File that maps gene names from CMOB datasets to their corresponding protein IDs in the STRING. This mapping allows researchers to easily translate gene-centric CMOB data into the protein-centric format required by STRING.
(2) CMOB offers Clinical Annotation, that is, EHR data, for patients who provided samples in the datasets. This resource supports the conducting of comprehensive phenotypic analyses using EHRs. The files in the auxiliary datasets are stored in a standardized format, allowing easy access and integration into analysis pipelines. Users can retrieve and read these files using simple scripts.
(3) CMOB also offers a suite of downstream Biological Attribute Analysis Tools for its omics datasets. These tools include gene differential expression (GDE) Rapaport et al. (2013) analysis at the single-gene level and pathway analysis Khatri et al. (2012) at the gene function set level. These resources aim to help researchers conduct classic and reliable biological validations of their data analysis frameworks.
More details on complementary resources can be found in the Supplementary Material Section E.
3.5 Datasets Release and Ethics Claims
Resources. All datasets and resources can be accessible via our GitHub repository (https://github.com/chenzRG/Cancer-Multi-Omics-Benchmark). We provide comprehensive guidelines for utilization. All files are ready for direct loading and analysis using standard Python data packages like Numpy or Pandas. We hope it can lower the barriers to entry for machine learning researchers interested in developing methods for cancer multi-omics data analysis, thereby encouraging rapid progress in the field.
Ethics Claims. As the original data source for CMOB, TCGA has established an Ethics, Law, and Policy Group to address critical ethical, legal, and social issues faced by researchers and participants. The group established informed consent guidelines for the effective and fair use of cancer genomic information, such as the "Data Use Certification Agreement", the "Human Subjects Protection and Data Access Policies", and the "Suggested Informed Consent Language for Prospective Collections".
4 Experiments on Selected Datasets
We show the experiments and results for selected tasks. We also provide an example of using the auxiliary dataset to integrate the CMOB dataset with STRING API to enable graph-based learning. We conducted all experiments on a server with an NVIDIA GeForce RTX 3090Ti GPU. The source codes and implementation details can be found in our GitHub.
4.1 Cancer Patient Classification
Experimental setup. We tested nine baseline cancer patient classification methods on four (one pan-cancer and three cancer subtype) patient classification datasets. We used the original feature scale data. The resulting labels from each method are compared with the true label for evaluation.
Results. Table 2 shows the cancer patient classification results. Overall, deep-based methods (Subtype-GAN, DCAP, MAUI, XOmiVAE, and MCluster-VAEs) have performed better than traditional statistical methods (SNF, NEMO, CIMLR, iClusterBayes, and moCluster). The performance of traditional methods is relatively uniform, and there is a significant gap between the best methods. This result shows the great potential of deep methods in cancer patient classification tasks.
Table 2: We tested nine baseline cancer patient classification methods on four patient classification datasets. The results are reported as PREC, NMI, and ARI.
Method Pan-cancer GS-BRCA GS-COAD GS-GBM
PREC NMI ARI PREC NMI ARI PREC NMI ARI PREC NMI ARI
SNF .643 .543 .475 .644 .523 .426 .625 .534 .432 .625 .544 .470
NEMO .656 .464 .356 .542 .444 .333 .644 .454 .333 .634 .406 .316
CIMLR .665 .365 .344 .655 .332 .345 .631 .343 .344 .647 .344 .323
iClusterBayes .747 .534 .433 .646 .524 .428 .637 .582 .434 .662 .506 .432
moCluster .725 .553 .557 .636 .630 .655 .749 .546 .652 .755 .734 .564
Subtype-GAN .844 .774 .748 .873 .734 .643 .851 .685 .648 .837 .625 .640
DCAP .845 .745 .636 .852 .743 .733 .852 .667 .655 .825 .642 .522
MAUI .859 .758 .625 .844 .792 .742 .882 .635 .696 .874 .741 .691
XOmiVAE .894 .795 .774 .843 .753 .761 .923 .752 .732 .946 .791 .737
MCluster-VAEs .883 .776 .763 .852 .784 .766 .895 .743 .727 .913 .783 .718
Table 3: We conducted missing value imputation experiments on five types of transcriptomics data with three different missing rates (70%, 50%, 30%). The results are reported as RMSE and MAE.
Data Missing Mean KNN MICE SVD SPEC GRAPE GAIN
RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE
BRCA 70% .119 .092 .109 .081 .106 .079 .099 .076 .104 .076 .127 .099 .117 .089
50% .119 .092 .103 .075 .090 .066 .086 .063 .090 .063 .131 .101 .114 .087
30% .119 .092 .099 .075 .084 .062 .080 .058 .088 .058 .131 .102 .112 .085
COAD 70% .101 .077 .099 .073 .093 .068 .089 .067 .094 .069 .102 .077 .104 .079
50% .101 .077 .091 .066 .079 .058 .077 .057 .076 .055 .110 .075 .103 .079
30% .102 .077 .086 .063 .076 .056 .072 .053 .071 .051 .105 .070 .103 .078
GBM 70% .122 .096 .106 .080 .097 .073 .096 .074 .110 .084 .125 .117 .122 .095
50% .122 .096 .097 .073 .084 .063 .082 .063 .084 .061 .145 .116 .115 .089
30% .122 .096 .093 .070 .080 .060 .078 .062 .083 .058 .146 .117 .114 .088
LGG 70% .131 .104 .109 .083 .095 .072 .097 .074 .153 .124 .152 .123 .132 .095
50% .131 .103 .098 .074 .082 .061 .081 .061 .082 .062 .151 .123 .129 .102
30% .131 .103 .094 .071 .078 .058 .076 .057 .074 .056 .151 .123 .123 .097
OV 70% .124 .098 .122 .094 .118 .091 .112 .088 .161 .130 .127 .101 .126 .099
50% .124 .098 .109 .083 .102 .078 .100 .075 .098 .078 .126 .099 .125 .098
30% .124 .098 .103 .078 .098 .075 .093 .071 .090 .069 .126 .099 .124 .097
Figure 2: The experiment results of (1) the external API example usage in three datasets and the downstream analysis on (2) pathway and (3) GDE in GS-BRCA.
4.2 Omics Data Imputation
Experimental setup. We tested six baseline missing value imputation methods on five omics data imputation datasets (mRNA). Values were independently and randomly removed for each feature before the imputation under three different missing rates (30%, 50%, 70%). The imputation methods receive the visible values, which are not removed as input, to predict the missing values.
Results. Table 3 shows the missing data imputation results. Overall, matrix decomposition methods (SVD, Spectral) have performed better than deep learning-based methods (GAN, GNN). This suggests that these methods might have captured inherent properties and the low-rank nature of the data, demonstrating the potential of matrix decomposition methods in imputation. This result indicates that traditional matrix decomposition methods still perform well in predicting missing mRNA values, whereas deep learning-based methods have room for improvement.
4.3 Integration Example of External Resources
Experimental setup. Here, we show an example of using the complementary resources to extend CMOB to an external API, i.e., STRING. STRING offers PPI networks defined by domain knowledge of molecular biology. These networks enable the organization of CMOB genetic profiles into a graph structure. Here, we employ SHapley Additive exPlanations (SHAP) Zhuang et al. (2023) and Graph Convolutional Networks on Multi-omics (GCNM) Li et al. (2021b), which are ConvGNN-based learning models for patient identification tasks. SHAP learns the gene feature from the CMOB pan-cancer dataset based on the gene node-connection information from the STRING PPI network for classifying samples.
Results. Table (1) in Figure 2 shows the external API experiment results. Overall, we successfully integrate CMOB omics data with a general PPI network from the STRING database and perform pan-cancer sample classification. The graph-based method can make better use of prior knowledge and shows improved performance. Based on the results of SHAP on GS-BRCA datasets, we conduct downstream analysis using the biological attribute analysis tools in complementary resources as an example. Figure (2) and (3) in Figure 2 show the results of GDE and pathway analysis. Overall, the results prove the biological divergence of the classification decisions. More detailed explanations of analysis results can be found in the Supplementary Material Section E.
5 Conclusion
Machine learning has shown great potential in cancer multi-omics analysis and advancing precision medicine. However, the most established and well-used public multi-omics data resource, TCGA, presents usability challenges, particularly for researchers without a biomedical background. To address this gap, our Cancer Multi-Omics Benchmark (CMOB) is the first large-scale cancer multi-omics benchmark: CMOB provides 20 cancer multi-omics datasets covering 32 cancers and complementary resources accessible and usable for machine learning researchers without significant preparation and expertise. We conduct extensive experiments on selected datasets to evaluate suitable machine learning baselines for specific tasks. Through CMOB, we aim to support the further catalyzing algorithmic and scientific innovation in cancer multi-omics data analysis.
Appendix A Key Information about CMOB
A.1 CMOB Structure
Here, we present the organizational structure of the CMOB, detailing its main components and resources. The CMOB repository is structured into three primary sections: Main Datasets, Baseline Models and Metrics, and Complementary Resources.
In Main Datasets, the repository hosts a comprehensive collection of tasks-ready cancer multi-omics datasets, stored primarily as .csv (Comma-Separated Values) files.
In Baseline Models and Metrics, the repository provides source codes of baseline models and evaluation metrics for different tasks, typically implemented in Python or R code.
In Complementary Resources, the repository encompasses additional tools and resources that complement the main datasets and downstream omics analysis needs, implemented as .csv or .py files.
CMOB
├── Main Datasets
│   ├── Pan-Cancer Dataset [Labeled]
│   │   └── Pan-Cancer
│   │       └── Aligned
│   │           └── mRNA / miRNA / CNV / Methy / Label.csv
│   ├── Cancer Subtype Datasets [Unlabeled]
│   │   ├── ACC
│   │   ├── KIRP
│   │   ├── KIRC
│   │   ├── LIHC
│   │   ├── LUAD
│   │   ├── LUSC
│   │   ├── PRAD
│   │   ├── THCA
│   │   └── THYM
│   │   (ABOVE)└── Original & Aligned & Top
│   │              └── mRNA / miRNA / CNV / Methy.csv
│   ├── Golden-Standard Cancer Subtype Datasets [Labeled]
│   │   ├── GS-COAD
│   │   ├── GS-BRCA
│   │   ├── GS-GBM
│   │   ├── GS-LGG
│   │   └── GS-OV
│   │   (ABOVE)└── Original & Aligned & Top
│   │              └── mRNA / miRNA / CNV / Methy / Label.csv
│   ├── Omics Data Imputation Datasets [Unlabeled]
│   │   ├── Imp-COAD
│   │   ├── Imp-BRCA
│   │   ├── Imp-GBM
│   │   ├── Imp-LGG
│   │   └── Imp-OV
│   │   (ABOVE)└── Top
│   │              └── mRNA / miRNA / CNV / Methy.csv
├── Baseline Models and Metrics
│   ├── Classification and Clustering Tasks
│   │   ├── Baselines.py/r
│   │   └── Metrics.py
│   └── Omics Data Imputation Tasks
│       ├── Baselines.py
│       └── Metrics.py
└── Complementary Resources
    ├── Mapping File
    │   └── CMOB_STRING_mapping.csv
    ├── Clinical Annotation
    │   └── Clinical_Rec.csv
    └── Biological Attribute Analysis Tools
        └── Analysis_Tools.py
A.2 Dataset Format
CMOB uses .csv files to manage and store all omics datasets, widely favored in biomedical research, including multi-omics studies. .csv files are plain-text files where data is separated by commas. They maintain a straightforward structure, with rows representing individual data records and columns representing different attributes or variables. Despite its simplicity, .csv remains efficient even with large datasets encountered in genomic and proteomic studies.
For example, consider a simplified .csv file containing mRNA data for a set of gene features (rows) across several patient samples (columns):
Feature Sample1 Sample2 Sample3 Sample4
GeneA 0.23 0.18 0.35 0.21
GeneB 0.56 0.49 0.52 0.58
GeneC 0.19 0.22 0.15 0.17
GeneD 0.08 0.10 0.09 0.12
In this example:
  • Each row corresponds to a specific gene (GeneA, GeneB, GeneC, GeneD).
  • Each column represents a different sample (Sample1, Sample2, Sample3, Sample4).
  • The numeric values in the cells denote the expression levels of each gene in each sample.
Both Python and R provide built-in functions and libraries to read, write, and analyze .csv files.
A.3 Recruited Cancer
The CMOB database contains multi-omics data for 32 types of cancer. The full names and abbreviations as shown in the following table:
Table 4: Cancer types and abbreviations in CMOB
No. Full Name Abbreviation
1 Acute Myeloid Leukemia LAML
2 Adrenocortical Cancer ACC
3 Bladder Urothelial Carcinoma BLCA
4 Brain Lower Grade Glioma LGG
5 Breast Invasive Carcinoma BRCA
6 Cervical & Endocervical Cancer CESC
7 Cholangiocarcinoma CHOL
8 Colon Adenocarcinoma COAD
9 Diffuse Large B-cell Lymphoma DLBC
10 Esophageal Carcinoma ESCA
11 Head & Neck Squamous Cell Carcinoma HNSC
12 Kidney Chromophobe KICH
13 Kidney Clear Cell Carcinoma KIRC
14 Kidney Papillary Cell Carcinoma KIRP
15 Liver Hepatocellular Carcinoma LIHC
16 Lung Adenocarcinoma LUAD
17 Lung Squamous Cell Carcinoma LUSC
18 Mesothelioma MESO
19 Ovarian Serous Cystadenocarcinoma OV
20 Pancreatic Adenocarcinoma PAAD
21 Pheochromocytoma & Paraganglioma PCPG
22 Prostate Adenocarcinoma PRAD
23 Rectum Adenocarcinoma READ
24 Sarcoma SARC
25 Skin Cutaneous Melanoma SKCM
26 Stomach Adenocarcinoma STAD
27 Testicular Germ Cell Tumor TGCT
28 Thymoma THYM
29 Thyroid Carcinoma THCA
30 Uterine Carcinosarcoma UCS
31 Uterine Corpus Endometrioid Carcinoma UCEC
32 Uveal Melanoma UVM
A.4 Recruited Omics
CMOB recruited four types of omics data:
  • mRNA (mRNA expression) measures the levels of messenger RNA transcribed from genes, reflecting the active transcription of genetic information;
  • miRNA (miRNA expression) quantifies the levels of microRNAs and small non-coding RNA molecules. It is crucial for post-transcriptional regulation in gene expression;
  • Methy (DNA methylation) measures the addition of methyl groups to DNA, typically at cytosine bases. It influences gene expression by altering the DNA accessibility to transcriptional machinery.
  • CNV (copy number variations) represents variations in the number of copies of particular DNA segments. This omics affects gene dosage and contributes to cancer susceptibility.
Appendix B CMOB Pipelines
B.1 Recruited Omics
mRNA (mRNA expression) measures the levels of messenger RNA transcribed from genes, reflecting the active transcription of genetic information;
miRNA (miRNA expression) quantifies the levels of microRNAs and small non-coding RNA molecules. It is crucial for post-transcriptional regulation in gene expression;
Methy (DNA methylation) measures the addition of methyl groups to DNA, typically at cytosine bases. It influences gene expression by altering the DNA accessibility to transcriptional machinery.
CNV (copy number variations) represents variations in the number of copies of particular DNA segments. This omics affects gene dosage and contributes to cancer susceptibility.
B.2 Omics Data Processing Pipelines
Here are the details for processing different omics:
(i) For transcriptomics (mRNA and miRNA) data:
  1. 1.
    STEP 1: Identify Transcriptomics Data Trace the data by "experimental_strategy" in the metadata, marked as "mRNA-Seq" or "miRNA-Seq". Check if "data_category" is marked as "Transcriptome Profiling".
  2. 2.
    STEP 2: Determine Experimental Platform Identify the experimental platform from metadata, such as "platform: Illumina" or "workflow_type: BCGSC miRNA Profiling".
  3. 3.
    STEP 3: Convert Gene-Level Estimates For data from the Hi-Seq platform like Illumina, use the R package edgeR Robinson et al. (2010) to convert the scaled estimates in the original gene-level RSEM to FPKM.
  4. 4.
    STEP 4: Filter Non-Human miRNA For "miRNA-Seq" data from Illumina GA and Agilent array platforms, identify and remove non-human miRNA expression features using species annotation from databases like miRBase Kozomara et al. (2019).
  5. 5.
    STEP 5: Eliminate Noise Identify and eliminate features with zero expression levels in more than 10% of samples or missing values (designated as N/A).
  6. 6.
    STEP 6: Apply Logarithmic Transformation Apply a logarithmic transformation to get the log-converted mRNA and miRNA data.
(ii) For genomic (CNV) data:
  1. 1.
    STEP 1: Identify CNV Alterations in Metadata Examine how alterations in gene copy-number are recorded in metadata using key descriptions like "Calls made after normal contamination correction and CNV removal using thresholds."
  2. 2.
    STEP 2: Filter Somatic Mutations Use keyword filtering to capture only somatic mutations, excluding germline mutations by retaining only those marked as ’somatic."
  3. 3.
    STEP 3: Identify Recurrent Alterations Use the R package GAIA al. SMe (2021) to identify recurrent alterations in the cancer genome from raw data that denote all aberrant regions resulting from copy number variation segmentation.
  4. 4.
    STEP 4: Annotate Genomic Regions Use the R package BiomaRt Durinck et al. (2005) to annotate the aberrant recurrent genomic regions.
  5. 5.
    STEP 5: Save Annotated CNV Data Save the annotated results to get CNV data of significantly amplified or deleted genes.
(iii) For epigenomic (Methy) data:
  1. 1.
    STEP 1: Identify Methylation Regions in Metadata Examine how methylation is defined in metadata to map methylation regions to genes, using key descriptions like "Average methylation (beta-values) of promoters defined as 500bp upstream & 50 downstream of Transcription Start Site (TSS)" or "With coverage >= 20 in 70% of the tumor samples and 70% of the normal samples."
  2. 2.
    STEP 2: Normalize Methylation Data Implement a median-centering normalization to account for systematic biases and technical variations across samples using the R package limmaRitchie et al. (2015).
  3. 3.
    STEP 3: Select Promoters with Minimum Methylation For genes with multiple promoters, select the promoter with minimum methylation in the normal tissues.
  4. 4.
    STEP 4: Save Mapped Methylation Data Save the mapped value data where each entry corresponds to a specific gene or genomic region, along with corresponding methylation measurements.
Appendix C CMOB Tasks
C.1 Pan-cancer Classification
Let $X^{O}=\{x_{1},x_{2},...,x_{m}\}$ represent the multi-omics dataset, where each $x_{i}$ is a vector of features in $O$-th omics for the $i$-th sample. Let $Y$ denote the set of possible cancer types. The goal of cancer classification using multi-omics data is to predict the true label $y_{i}$ for each sample $x_{i}$ in $X$, where $y_{i}$ belongs to the set of possible cancer types $Y$. Cancer classification can be formulated as a supervised learning problem, where the objective is to learn a mapping function $f:X\rightarrow Y$ that accurately predicts the true labels for unseen samples based on their omics features.
C.2 Cancer Subtype Clustering
Cancer subtyping means categorizing patients into subgroups that exhibit differences in various aspects based on their multi-omics data. However, for most cancer types, especially rare cancers, the cancer subtyping tasks are still open questions under discussion. Thus, cancer subtyping tasks are typically clustering tasks without ground true labels. Let $X^{O}=\{x_{1},x_{2},...,x_{m}\}$ represent the multi-omics dataset, where each $x_{i}$ is a vector of features in $O$-th omics for the $i$-th sample. Let $k$ denote the set of possible cancer subtypes. The goal of cancer subtyping using multi-omics data is to assign each sample $x_{i}$ in $X$ into $k$ clusters $C=\{C_{1},C_{2},...,C_{k}\}$, such that each cluster $C_{i}$ represents a distinct cancer subtype based on the information from multiple omics data sources.
C.3 Golden-standard Subtype Classification
The cancer research community has thoroughly analyzed the subtypes of some of the most common cancer types in a previous study. Therefore, we consider these subtypes to be the true labels. The definition of golden-standard subtype identification is similar to the above Pan-cancer identification tasks. Golden-standard subtype identification task aims to assign each sample $x$ in the sample set $X$ to a cancer subtype $y$ in the set of all subtypes $Y$.
C.4 Omics Data Imputation
Let $X$ denote the original omics data with $m$ samples and $n$ features, represented as a matrix where $X_{ij}$ represents the value of the $i$-th sample for the $j$-th feature. Let $M$ denote the binary mask matrix of the same dimensions as $X$, where $M_{ij}=1$ if the value of $X_{ij}$ is observed (not missing), and $M_{ij}=0$ if it is missing. The goal of the imputation task is to estimate the missing values in $X$, denoted as $\hat{X}$, using the observed values and potentially additional information. Imputation can be formulated as $\hat{X}=f(X,M)$, where $f$ is the imputation function that takes as input the original omics data $X$ and the mask matrix $M$, and outputs the imputed matrix $\hat{X}$.
Appendix D Evaluation Metrics
D.1 Silhouette coefficient (SIL)
The silhouette coefficient measures the similarity between a sample and its classified subtype compared to the samples in the other subtypes to determine how appropriately samples in a dataset have been clustered. For a sample $i$, let $a(i)$ be the average distance from sample $i$ to other samples in the same cluster, and let $b(i)$ be the smallest average distance from sample $i$ to samples in a different cluster, minimized over clusters. The silhouette coefficient $SIL(i)$ for a sample $i$ is then defined as:
$SIL(i)=\frac{b(i)-a(i)}{\max\{a(i),b(i)\}}$
The silhouette coefficient ranges from -1 to 1, where a high value indicates that the sample is well-matched to its own cluster and poorly matched to neighboring clusters.
D.2 P-value of the log-rank test on survival time (LPS)
The log-rank test on survival time is a hypothesis test used to compare the survival distributions of two or more groups. The test statistic $X^{2}$ is calculated from the observed and expected number of events in each group over time. The p-value is then calculated from the test statistic under the null hypothesis that there is no difference in survival distributions between the groups. The LPS gives the log-transformed p-values of the log-rank test. It is calculated as below using the chi-square distribution with $k-1$ degrees of freedom:
$LPS=P(X^{2}\geq X^{2}_{observed})$
where $k$ is the number of groups being compared, $X^{2}_{observed}$ is the observed test statistic calculated from the data.
D.3 Precision (PREC)
Precision measures the accuracy of the positive predictions made by a classification or clustering model. It is defined as the ratio of true positive (TP) predictions to the total number of positive predictions made by the model:
$PREC=\frac{TP}{TP+FP}$
where $TP$ is the number of true positive predictions (instances correctly classified as positive), and $FP$ is the number of false positive predictions (instances incorrectly classified as positive).
D.4 Normalized mutual information (NMI)
Normalized mutual information measures the similarity between two clusterings of the same dataset. It measures the mutual dependence between the clustering result and the ground truth labels, normalized by the average entropy of the two clusterings. Let $C$ be the clustering result and $G$ be the ground truth labels. Then, NMI is calculated as:
$NMI(C,G)=\frac{I(C,G)}{\sqrt{H(C)\cdot H(G)}}$
where $I(C,G)$ is the mutual information between $C$ and $G$, $H(C)$ and $H(G)$ are the entropies of $C$ and $G$, respectively.
D.5 Adjusted rand index (ARI)
Adjusted rand index measures the similarity between two clusterings of the same dataset. It measures the agreement between the pairs of samples assigned to the same or different clusters in the two compared clusterings, adjusted for chance. ARI is calculated as:
$ARI(C,G)=\frac{{a+b}}{{\binom{n}{2}}}-\frac{{a\cdot(a-1)+b\cdot(b-1)}}{{\binom{n}{2}}}$
where $a$ is the number of pairs of samples that are in the same cluster in both $C$ and $G$, $b$ is the number of pairs of samples that are in different clusters in both $C$ and $G$, $n$ is the total number of samples, and $\binom{n}{2}$ is the number of all possible pairs of samples.
D.6 Mean absolute error (MAE)
Mean absolute error measures the average absolute difference between the imputed values and the true values as below:
$MAE=\frac{1}{n}\sum_{i=1}^{n}|\hat{Y}_{i}-Y_{i}|$
where $n$ is the number of imputed values, $\hat{Y}_{i}$ is the imputed value for observation $i$ and $Y_{i}$ is the true value for observation $i$.
D.7 Root mean squared error (RMSE)
Root mean squared error measures the square root of the average squared difference between the imputed values and the true values as below:
$RMSE=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(\hat{Y}_{i}-Y_{i})^{2}}$
where $n$ is the number of imputed values, $\hat{Y}_{i}$ is the imputed value for observation $i$ and $Y_{i}$ is the true value for observation $i$.
Table: The evaluation metrics and number of baselines used for all tasks in CMOB. Tasks #Baselines Metrics Pan-cancer Classification 10 PREC, NMI, ARI Cancer Subtype Clustering 10 SIL, LPS Golden-standard Subtype Classification 10 PREC, NMI, ARI Omics Data Imputation 7 MAE, RMSE
Appendix E Complementary Resources
E.1 Differential Gene Expression Analysis.
Differential gene expression analysis has been a cornerstone of transcriptomic studies. In this analysis, we compare gene expression levels between different experimental conditions or sample groups to identify genes that are significantly upregulated or downregulated. Statistical tests such as t-tests or non-parametric tests are commonly used for this purpose. For example, comparing gene expression profiles between cancer patients and healthy controls to identify genes that are dysregulated in cancer. Genes with significant differences in expression levels may be further investigated as potential biomarkers or therapeutic targets. For example, researchers performed differential gene expression analysis on RNA-seq data from Alzheimer’s disease patients and healthy controls. This analysis identified a panel of differentially expressed genes implicated in neuroinflammation and synaptic dysfunction, showing molecular pathways associated with Alzheimer’s disease progression.
We calculated the log2 fold change in gene abundance between pairwise groups and determined the significance of expression changes using Student’s t-test. P-values were adjusted using the Benjamini-Hochberg procedure to correct the false discovery rate. We considered a gene to be significant if it had an adjusted p-value less than 0.05 and a log2 fold change greater than or equal to 1.2. Based on their fold changes, the resulting DEGs were categorized into up-regulated and down-regulated sets and can be utilized for subsequent analysis phases.
Among the identified DEGs, several genes have been extensively reported as being associated with cancer progression. Notable examples include BRCA1, WNT4, and NOTCH2. BRCA1 is well-known for its involvement in hereditary breast cancer and plays essential roles in cell cycle regulation, DNA damage response, and transcriptional control Miki et al. (1994). Dysregulation of the WNT4 gene, which encodes a protein belonging to the Wnt signaling pathway, has been linked to tumor growth, invasion, and metastasis Peradziryi et al. (2011). Similarly, the NOTCH2 gene, a member of the Notch receptor family, is critical in cell fate determination, development, and tissue homeostasis and has been implicated in tumor initiation, progression, and therapy resistance Gridley (2003).
E.2 Pathway Analysis.
Pathway analysis is a critical step in interpreting the biological significance of DEGs. By mapping DEGs to known biological pathways, researchers can gain insights into the underlying mechanisms and potential functional impacts of gene expression changes. In CMOB, pathway analysis is performed using established databases such as KEGG. These databases provide curated information on metabolic pathways, signaling cascades, and gene ontologies. DEGs are input into pathway analysis tools to conduct the analysis, which then identifies overrepresented pathways among the upregulated and downregulated gene sets.
For instance, pathway enrichment analysis might reveal that upregulated DEGs in cancer samples are significantly associated with pathways involved in cell cycle regulation and apoptosis, while downregulated DEGs are linked to immune response pathways. Such findings can help to identify potential therapeutic targets and elucidate the molecular basis of disease.
In our approach, we utilize Fisher’s exact test or hypergeometric test to evaluate the significance of pathway enrichment. Adjustments for multiple testing are performed using the Benjamini-Hochberg procedure, with pathways considered significant at an adjusted p-value threshold of less than 0.05. The pathway analysis results are visualized using enrichment plots and pathway diagrams, which highlight key genes and interactions within the enriched pathways.
Appendix F Data Source Ethics and Policies
The ultimate goal of data source ethics and policies was to develop research policies maximizing public benefit from the data that were by these ethical and legal guidelines, ensuring: (1) Protection of human participants in the project, including their privacy; (2) Secure and compliant access to TCGA data; (3) Timely data release to the research community; (4) Initial scientific publication by the data producers; (5) These policies have influenced the field of cancer genomics and will continue to serve as a guide for future genomic research projects.
F.1 Human Subjects Protection and Data Access Policies:
NCI and NHGRI developed a set of policies to protect the privacy of participants donating specimens to TCGA. TCGA’s informed consent policy, data access policy, and information about compliance with the HIPAA Privacy Rule are included.
F.2 Data Use Certification Agreement:
Researchers must agree to A set of policies before gaining access to TCGA data. This agreement ensures that researchers pursuing a research question requiring controlled-access data comply with TCGA policies, such as maintaining participants’ privacy, securely accessing the data, and following TCGA publication guidelines.
F.3 Suggested Informed Consent Language for Prospective Collections:
An example informed consent document that TCGA suggested Tissue Source Sites use when collecting specimens from prospective project participants. This document helps ensure that patients considering donating tissue specimens to human genomics research programs such as TCGA recognize the risks and benefits of participation and understand the nature of their inclusion in the project.
F.4 Sharing Data from Large-scale Biological Research Projects:
Principles for sharing and publishing genomic data to maximize public benefit developed at a meeting in Fort Lauderdale sponsored by the Wellcome Trust. These “Fort Lauderdale Principles” informed the original TCGA publication guidelines, which balance making genomic data immediately available for research use with protecting the original owner’s initial publication rights.
F.5 Considerations for Open Release of Genomic Data from Human Cancer Cell Lines:
An explanation of the factors considered in the decision by NCI and NHGRI to release genomic data and information from the Cancer Cell Line Encyclopedia as open-access data.
Appendix G Limitations & Broader Impact
This research proposes a benchmark for cancer multi-omics data analysis. However, we collected all data from TCGA sources and did not conduct wet experiments to introduce new data further. Consequently, the data is limited and influenced by the specific cohorts and methodologies used in TCGA, which may not fully represent the diversity of cancer types or the broader patient population. Cancer omics data also raises ethical issues, particularly in cancer risk prediction and the development of anti-cancer drugs, which could have potentially harmful or controversial functions. The use of omics data for predictive purposes can lead to concerns about privacy, discrimination, and the psychological impact on individuals who are identified as high-risk. Additionally, designing drugs based on omics data can lead to unintended side effects and ecological impacts if not carefully regulated.
Nevertheless, we believe that omics data has great potential to benefit society. It can lead to more personalized and effective treatments, early cancer detection, and a better understanding of cancer biology. Negative impacts can be mitigated through stringent industry regulations, ethical guidelines, and legislation to ensure responsible use and data protection. The proposed benchmark helps the community develop new cancer omics data analysis algorithms and evaluate the performance of existing models. By providing a standardized framework, we aim to facilitate advancements in cancer research and improve the reproducibility and comparability of different computational approaches.